# Fast Penalized Generalized Estimating Equations for Large Longitudinal Functional Datasets

### v1

@@ Abstract
Longitudinal binary or count functional data are common in neuroscience, but are often too large to analyze with existing functional regression methods. We propose one-step penalized generalized estimating equations that supports continuous, count, or binary functional outcomes and is fast even when datasets have a large number of clusters and large cluster sizes. The method applies to both functional and scalar covariates, and the one-step estimation framework enables efficient smoothing parameter selection, bootstrapping, and joint confidence interval construction. Importantly, this semi-parametric approach yields coefficient confidence intervals that are provably valid asymptotically even under working correlation misspecification. By developing a general theory for adaptive one-step M-estimation, we prove that the coefficient estimates are asymptotically normal and as efficient as the fully-iterated estimator; we verify these theoretical properties in extensive simulations. Finally, we apply our method to a calcium imaging dataset published in Nature, and show that it reveals important timing effects obscured in previous non-functional analyses. In doing so, we demonstrate scaling to common neuroscience dataset sizes: the one-step estimator fits to a dataset with 150,000 (binary) functional outcomes, each observed at 120 functional domain points, in only $\sim 13.5$ minutes on a laptop without parallelization. We release our implementation in the fastFGEE package.
@@
Keywords: functional data analysis, longitudinal data analysis, calcium imaging, generalized estimating equations, one-step estimators
Introduction
@@
Neuroscience studies in animal models provide an invaluable tool to identify the neural mechanisms underpinning brain function and their relationship with psychiatric disorders. Researchers can estimate moment-by-moment associations between experimental covariates (e.g. behavior) and the activity of hundreds of neurons per animal, with widely-used in vivo brain recording techniques like calcium imaging (Grienberger et al.,, 2022) and Neuropixels (Jun et al.,, 2017). A neuroscientist might study brain-behavior associations on, for example, a learning task in which an animal learns to press a lever for a food reward. These tasks are often performed over hundreds of experimental replicates called “trials” (longitudinal observations akin to “patient visits”). Each trial might be defined as a five second interval starting at extension of the lever and ending with delivery of the food reward. To test whether, for example, mean neural activity is higher on trials when animals press the lever, a common strategy is to analyze scalar summaries of each trial’s neuronal firing activity. For instance, analysts might calculate a firing rate of neuron $i$ on trial $j$ by averaging the response, $Y_{i,j}(s)\in\{0,1\}$, across within-trial timepoints indexed by $s$: $\bar{Y}_{i,j}=\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}Y_{i,j}(s)$, where $\mathcal{S}\subset[0,5]$ denotes a grid of timepoints at which the outcome is observed. One might then test whether $\mathbb{E}[\bar{Y}_{i,j}\mid X_{i,j}=1]-\mathbb{E}[\bar{Y}_{i,j}\mid X_{i,j}=0]\neq 0$, where $X_{i,j}$ is an indicator that the animal on which neuron $i$ was recorded pressed the lever on trial $j$. This analysis approach is parsimonious but discards important temporal information by summarizing across trial timepoints, $s$.
@@
Alternatively, the neural response of each five second trial can be conceptualized as a functional outcome, with within-trial timepoints, $s$, representing locations along the functional domain. This allows one to apply functional data analysis (FDA, Crainiceanu et al., (2024)) techniques to test how brain–behavior relationships evolve within and across trials (Loewinger et al.,, 2025). FDA offers a flexible framework to analyze a wide range of neuroscience studies, as it allows both the responses and the covariates (e.g. behavior) to be functional. The size and complexity of our dataset, however, require specialized FDA methods. First, analyses must account for the longitudinal structure, since each neuron’s activity is collected across many trials. Second, the large number of clusters (i.e. neurons) and large cluster sizes (i.e. number of trials recorded for a given neuron), often in the hundreds or thousands, make many longitudinal FDA methods for binary or count functional outcomes computationally impractical.
@@
We provide an implementation that supports functional data observed on regular, irregular, dense and sparse grids with functional and/or scalar covariates (see Appendix A.4 for implementation details and Section 7 for links to our Github repos). We further propose adoption of working correlation structures that allow for use of algorithms to efficiently construct and invert large working covariance matrices. We also propose fast strategies for smoothing parameter tuning, cluster bootstrapping, and joint confidence interval construction. We review the literature in Section 1, present our estimator in Section 2, provide theoretical results in Section 3, simulations in Section 4, and a data application in Section 5.
1 Related Literature
We focus on the longitudinal function-on-scalar regression literature, where a wide range of conditional and marginal methods have been proposed (Eckardt et al.,, 2024; Sergazinov et al.,, 2023; Zhu et al.,, 2019; Scheipl et al.,, 2016; Shou et al.,, 2015; Brockhaus et al.,, 2015; Scheipl et al.,, 2015; Zipunnikov et al.,, 2014; Greven et al.,, 2011; Morris and Carroll,, 2006; Guo,, 2002).
@@ 1 Related Literature
Functional mixed models are a versatile conditional strategy for longitudinal FDA. For non-Gaussian functional outcomes, many existing approaches do not scale well to large cluster sizes or cluster numbers; see discussion and comparison in Cui et al., (2022). Cui et al., (2022); Loewinger et al., (2025); Zhou et al., (2025) proposed a fast functional mixed models approach based on univariate mixed models fits at each functional domain point. These rely on a cluster bootstrap for inference, however, which can be slow for large datasets. Moreover, for non-Gaussian outcomes these approaches yield coefficient estimates that are only interpretable as conditional on the random effects. In many applications, estimates with marginal interpretations are desirable.
@@
Functional GEE and Quadratic Inference Functions (QIF) are marginal methods for longitudinal function-on-scalar regression. Qu and Li, (2005) and Guha Niyogi and Zhong, (2025) proposed QIF-based methods applicable to FDA, but, to the best of our understanding, these works focused on a single observation of a functional outcome per subject. Chen et al., (2013) proposed a penalized GEE for longitudinal FDA that serves as part of the inspiration for our work. The method, however, requires inverting an $n_{i}L\times n_{i}L$ matrix at each step of model fitting, where $n_{i}$ is the size of cluster $i$, and $L$ is the number of points in the functional domain. Li et al., (2022) proposed a marginal estimator for continuous outcomes, but as we show in simulations, it does not scale to large cluster sizes, and has not been extended to binary or count outcomes. Taken together, marginal approaches for longitudinal functional regression with binary or count outcomes do not scale well, thereby limiting their widespread adoption.
Finally, there is a rich literature on penalized longitudinal marginal models for high dimensional covariates with sparsity; see Xia and Shojaie, (2024) and references therein. However, we focus on functional data with low dimensional covariates where we do not encourage sparsity.
@@ 2 Methods
We begin by introducing notation, adopting that used in Li et al., (2022) where possible. We suppose that we observe the functional outcome $Y_{i,j}(s)$ at point $s$ for cluster $i\in[N]\coloneqq\{1,\ldots,N\}$, at longitudinal observation (e.g. trial or visit) $j\in[n_{i}]$. We express grids as regular (i.e. $n_{i}=n_{i}(s)~{}~{}\forall~{}s\in\{s_{1},\ldots,s_{L}\}$) and evenly spaced for ease of notation, but our methods also apply to irregular and unevenly spaced grids. We denote $\bm{Y}_{i}(s)\in\mathbb{R}^{n_{i}}$ as the functional outcome vector at point $s$ for cluster $i$, concatenating all observations $Y_{i,j}(s)$ for $j\in[n_{i}]$, and write $\bm{Y}_{i}=[\bm{Y}_{i}(s_{1})^{T},\ldots,\bm{Y}_{i}(s_{L})^{T}]^{T}\in\mathbb{R}^{n_{i}L}$. We denote covariate vector $\mathbf{X}_{i,j}\in\mathbb{R}^{q}$ for cluster $i$ on observation $j$, and $\mathbf{X}_{i}=[\mathbf{X}_{i,1},\ldots,\mathbf{X}_{i,n_{i}}]^{T}\in\mathbb{R}^{n_{i}\times q}$. We write covariates as scalar for ease of notation, though our method and theory applies to functional covariates.
@@ 2.1 Functional Generalized Estimating Equations
We consider the marginal function-on-scalar regression
$g(\mathbb{E}[Y_{i,j}(s)\mid\mathbf{X}_{i,j}])=\eta_{i,j}(s),~{}~{}\eta_{i,j}(s)=\beta_{0}(s)+\sum_{r=1}^{q}X_{i,j,r}\beta_{r}(s),~{}~{}\text{Cov}\left(\bm{Y}_{i}\mid\mathbf{X}_{i}\right)=\mathbb{V}_{i}^{*}$ (1)
where $g$ is a link function and $\beta_{r}(\cdot)$ is a (smooth) coefficient function for covariate $r\in[q]$. We let $\mu_{i,j}(s)=g^{-1}(\eta_{i,j}(s))$ denote the mean function. We now discuss estimation of $\mu_{i,j}(s)$ with spline basis expansions of the $\beta_{r}(\cdot)$, although our methods can be used for other basis functions. For example, denoting $\mathbf{B}(s)=[B_{1}(s),\ldots,B_{m}(s)]^{T}\in\mathbb{R}^{m}$ as a set of $m$ B-spline basis functions, we can represent the functional coefficients $\beta_{r}(s)=\sum_{d=1}^{m}\theta_{r,d}B_{d}(s)$. We denote $\bm{\theta}_{r}=[\theta_{r,1},\ldots,\theta_{r,m}]^{T}\in\mathbb{R}^{m}$ as an unknown parameter vector associated with covariate $r$, $\mathbf{B}=[\mathbf{B}(s_{1}),\ldots,\mathbf{B}(s_{L})]^{T}\in\mathbb{R}^{L\times m}$, and the linear predictor for a full observation of the functional outcome as $\bm{\eta}_{i,j}=[\eta_{i,j}(s_{1}),\ldots,\eta_{i,j}(s_{L})]^{T}=\mathbf{B}\bm{\theta}_{0}+\sum_{r=1}^{q}X_{i,j,r}\mathbf{B}\bm{\theta}_{r}$. We further define $\mathbb{X}_{i,j}=[\mathbf{B},~{}X_{i,j,1}\mathbf{B},\ldots,X_{i,j,q}\mathbf{B}]\in~{}\mathbb{R}^{L\times p}$, where $p=m(1+q)$. We then have that $\bm{\eta}_{i,j}=\mathbb{X}_{i,j}\bm{\theta}$, where $\bm{\theta}=[\bm{\theta}_{0}^{T},~{}\bm{\theta}_{1}^{T},~{}\ldots,~{}\bm{\theta}_{q}^{T}]^{T}\in\mathbb{R}^{p}$. Thus, we can estimate the functional coefficient vector, $\bm{\beta}_{r}=[\beta_{r}(s_{1}),~{}\ldots,~{}\beta_{r}(s_{L})]^{T}\in\mathbb{R}^{L}$, by estimating $\bm{\theta}$ and calculating $\widehat{\bm{\beta}}_{r}=\mathbf{B}\widehat{\bm{\theta}}_{r}$.
We semi-parametrically estimate the $\bm{\theta}$ with the penalized spline-based fGEE proposed in Chen et al., (2013). This assumes no likelihood and, if $\mu_{i,j}(s)$ is correctly specified, yields valid inference for $\{\bm{\beta}_{r}(s)\}_{s,r}$ even if $\text{Cov}(\bm{Y}_{i}\mid\mathbf{X}_{i})$ is misspecified. Specifically, the mean model parameters $\bm{\theta}$ are estimated as the root of the penalized estimating equation
@@
where $\mathbb{V}_{i}\in\mathbb{R}^{n_{i}L\times n_{i}L}$ is the working covariance matrix for cluster $i$ (whose true covariance matrix is $\mathbb{V}_{i}^{*}$), $\mathbb{D}_{i}=\frac{\partial\bm{\mu_{i}}(\bm{\theta})}{\partial\bm{\theta}}$, $\mathbb{X}_{i}=[\mathbb{X}_{i,1}^{T},\ldots,\mathbb{X}_{i,n_{i}}^{T}]^{T}\in\mathbb{R}^{n_{i}L\times p}$, and $\bm{\mu}_{i}=[\bm{\mu}_{i,1}^{T},\ldots,\bm{\mu}_{i,n_{i}}^{T}]^{T}\in\mathbb{R}^{n_{i}L}$. The pre-specified penalty matrix, $\mathbb{S}\in\mathbb{R}^{p\times p}$, is associated with the diagonal matrix of smoothing parameters $\Lambda\in\mathbb{R}^{p\times p}$. Although no likelihood is adopted, the estimating equation (2) can be derived from the score equations from, for example, an exponential dispersion family (Liang and Zeger,, 1986); we add the penalty term for improved estimation in finite samples. Compared to a working independence matrix $\mathbb{V}_{i}=I_{n_{i}L}$, the estimation of $\bm{\theta}$ can be made more efficient and accurate by exploiting correlation, in both functional and longitudinal directions, by choosing the working covariance matrix $\mathbb{V}_{i}$ to estimate $\mathbb{V}_{i}^{*}$. Although such choices for $\mathbb{V}_{i}$ in this fGEE model yield desirable statistical properties for longitudinal FDA, estimation is computationally intensive: estimating $\bm{\theta}$ based on equation (2) requires inversion of the $n_{i}L\times n_{i}L$ covariance matrix $\mathbb{V}_{i}$ for each cluster $i$, at each step in an optimization procedure.

### v2

@@ Abstract
Longitudinal binary or count functional data are common in neuroscience, but are often too large to analyze with existing functional regression methods. We propose one-step penalized generalized estimating equations that supports generalized functional outcomes (e.g., count, binary, proportion, continuous-valued) and is fast even when datasets have a large number of clusters and large cluster sizes. The method applies to functional and scalar covariates and the one-step estimation framework enables efficient smoothing parameter selection and joint confidence interval construction. Importantly, this semi-parametric approach yields coefficient confidence intervals that are provably valid asymptotically even under working correlation misspecification. By developing a general theory for adaptive one-step M-estimation, we prove that the coefficient estimates are asymptotically normal and as efficient as the fully-iterated estimator; we verify these theoretical properties in simulations. We illustrate the benefits of our approach for analyzing large-scale neural recordings by applying it to a recent calcium imaging dataset published in Nature. We show that our method reveals important timing effects obscured in non-functional analyses. In doing so, we also demonstrate scaling to common neuroscience dataset sizes: the one-step estimator fits to a dataset with 150,000 (binary) functional outcomes, each observed at 120 functional domain points, in only $\sim 6.5$ minutes on a laptop without parallelization. We release our methods in the R package fastfGEE, which supports a wide range of link functions and working covariances.
@@
Keywords: calcium imaging, functional data analysis, generalized estimating equations, longitudinal data analysis, one-step estimators
1 Introduction
@@
Neuroscience studies in animal models provide an invaluable tool to identify the neural mechanisms underpinning psychiatric disorders. Researchers can estimate moment-by-moment associations between experimental covariates (e.g. behavior) and the activity of hundreds of neurons per animal, with widely-used in vivo brain recording techniques like calcium imaging (Grienberger et al., 2022) and Neuropixels (Steinmetz et al., 2018). A neuroscientist might study brain-behavior associations on, for example, a learning task in which an animal learns to press a lever for a food reward. These tasks are often performed over hundreds of experimental replicates called “trials” (longitudinal observations akin to “patient visits”). Each trial might be defined as a five second interval starting at extension of the lever and ending with delivery of the food reward. To test whether, for example, mean neural activity is higher on trials when animals press the lever, a common strategy is to analyze scalar summaries of each trial’s neuronal firing activity. For instance, analysts might calculate a firing rate of neuron $i$ on trial $j$ by averaging the response, $Y_{i,j}(s)\in\{0,1\}$, across within-trial timepoints indexed by $s$: $\bar{Y}_{i,j}=\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}Y_{i,j}(s)$, where $\mathcal{S}\subset[0,5]$ denotes a grid of timepoints at which the outcome is observed. One might then test whether $\mathbb{E}(\bar{Y}_{i,j}\mid X_{i,j}=1)-\mathbb{E}(\bar{Y}_{i,j}\mid X_{i,j}=0)\neq 0$, where $X_{i,j}$ is an indicator that the animal on which neuron $i$ was recorded pressed the lever on trial $j$. This discards important temporal information, however, by summarizing across timepoints.
@@
Alternatively, the neural response of each five second trial can be conceptualized as a functional outcome, with within-trial timepoints, $s$, representing locations along the functional domain. This allows one to apply functional data analysis (FDA) techniques to test how brain–behavior relationships evolve within and across trials (Loewinger et al., 2025) as both the responses and the covariates (e.g., behavior) can be modeled as functional variables. The size and complexity of our dataset, however, require specialized FDA methods. First, analyses must account for the longitudinal correlation, since each neuron’s activity is collected across many trials. Second, the large number of clusters (500 neurons), large cluster sizes (300 trials/neuron), and densely sampled functional outcomes ($|\mathcal{S}|=120$ timepoints/trial) render many longitudinal methods for generalized functional outcomes computationally impractical. Our semiparametric approach models correlation in functional and longitudinal directions to improve statistical efficiency, and adopts a variance estimator robust to covariance misspecification to ensure valid inference.
@@
Our implementation supports functional data observed on regular, irregular, dense and sparse grids with functional and scalar covariates (see Appendix A.10). We propose fast strategies for smoothing parameter selection and joint confidence interval construction. Our general theory for adaptive one-step M-estimation may be of independent interest.
@@
Our methods build on the rich literature for conditional and marginal methods for longitudinal function-on-scalar regression (Zhu et al., 2019; Scheipl et al., 2016; Wang et al., 2016; Morris and Carroll, 2006; Ramsay and Silverman, 2005; Guo, 2002). Functional mixed models are a versatile conditional strategy for longitudinal FDA. For non-Gaussian functional outcomes, many existing approaches do not scale well to large cluster sizes or cluster numbers. For example, Functional Additive Mixed Models, implemented in refund::pffr() (Scheipl et al., 2016), can become computationally intensive when the number of clusters is in the low hundreds, even when specifying random functional intercept-only models: each participant adds $k_{b}$ columns to the design matrix, where $k_{b}$ is the number of spline bases used for the random intercept (Cui et al., 2022). Cui et al. (2022); Loewinger et al. (2025) proposed a fast functional mixed models approach based on univariate mixed models fits at each functional domain point. For non-Gaussian outcomes this relies, however, on a cluster bootstrap for inference, which can be slow for large datasets. Moreover, for non-Gaussian outcomes these approaches yield coefficient estimates that are only interpretable as conditional on the random effects. In many applications, estimates with marginal interpretations are desirable.
@@
Functional GEE and Quadratic Inference Functions (QIF) are marginal methods for longitudinal function-on-scalar regression. Qu and Li (2005) and Guha Niyogi and Zhong (2025) proposed QIF-based methods applicable to FDA, but, to the best of our understanding, these works focused on a single observation of a functional outcome per subject. Chen et al. (2013) proposed a penalized GEE for longitudinal FDA that serves as part of the inspiration for our work. The method, however, requires inverting an $n_{i}L\times n_{i}L$ matrix at each step of model fitting, where $n_{i}$ is the size of cluster $i$, and $L$ is the number of points in the functional domain. Li et al. (2022) proposed a marginal estimator for continuous data, but it has not been extended to generalized outcomes. Taken together, marginal approaches for longitudinal functional regression with generalized outcomes do not scale well, limiting their widespread adoption.
Our work builds on existing one-step approaches to scale GEE to large problem sizes. Cheng et al. (2013) establish that a one-step GEE cluster bootstrap yields asymptotically equivalent inference as a fully-iterated GEE cluster bootstrap (see Section 2.2). This provides a basis for the computational steps we use for cross-validation and confidence interval construction. Lipsitz et al. (2017) use a non-adaptive, unpenalized one-step estimator to scale non-functional GEE to large cluster sizes when adopting an exchangeable working correlation. By contrast, our theory and methods are for adaptive penalized one-step M-estimation and can be used for functional data with many working correlations. One-step estimators are also used to enable inference for GEE in high dimensional, sparse settings. For example, Xia and Shojaie (2022) propose a one-step estimator using a projected estimating equation to enable inference of linear functionals of high dimensional GEE coefficients in non-functional data settings. In contrast, we focus on functional data with low-dimensional covariates where we do not encourage sparsity.
Our paper is organized as follows. We present our methodology in Section 2, provide theoretical results in Section 3, simulations in Section 4, and a data application in Section 5.
@@ 2 Methods
We begin by introducing notation, adopting that used in Li et al. (2022) where possible. We suppose that we observe the functional outcome $Y_{i,j}(s)$ at point $s$ for cluster $i\in[N]\coloneqq\{1,\ldots,N\}$, at longitudinal observation (e.g. trial or visit) $j\in[n_{i}]$. We express grids as regular (i.e. $n_{i}=n_{i}(s)~~\forall~s\in\{s_{1},\ldots,s_{L}\}$) and evenly spaced for ease of notation, but our methods also apply to irregular and unevenly spaced grids. We denote $\bm{Y}_{i}(s)\in\mathbb{R}^{n_{i}}$ as the functional outcome vector at point $s$ for cluster $i$, concatenating all observations $Y_{i,j}(s)$ for $j\in[n_{i}]$, and write $\bm{Y}_{i}=[\bm{Y}_{i}(s_{1})^{T},\ldots,\bm{Y}_{i}(s_{L})^{T}]^{T}\in\mathbb{R}^{n_{i}L}$. We denote covariate vector $\mathbf{X}_{i,j}\in\mathbb{R}^{q}$ for cluster $i$ on observation $j$, and $\mathbf{X}_{i}=[\mathbf{X}_{i,1},\ldots,\mathbf{X}_{i,n_{i}}]^{T}\in\mathbb{R}^{n_{i}\times q}$. We write covariates as scalar for ease of notation, though our method and theory applies to functional covariates.
@@ 2.1 Functional Generalized Estimating Equations
We consider the marginal function-on-scalar regression with link function $g$,
$g\{\mathbb{E}(Y_{i,j}(s)\mid\mathbf{X}_{i,j})\}=\eta_{i,j}(s),~~\eta_{i,j}(s)=\beta_{0}(s)+\sum_{r=1}^{q}X_{i,j,r}\beta_{r}(s),~~\text{Cov}\left(\bm{Y}_{i}\mid\mathbf{X}_{i}\right)=\mathbb{V}_{i}^{*}$ (1)
where $\beta_{r}(\cdot)$ is a (smooth) coefficient function for covariate $r\in[q]$. We let $\mu_{i,j}(s)=g^{-1}(\eta_{i,j}(s))$ denote the (assumed) mean function; $\mathbb{V}_{i}^{*}$ represents the true within-cluster covariance, which is arbitrary and unknown. We now discuss estimation of $\mu_{i,j}(s)$ with spline basis expansions of the $\beta_{r}(\cdot)$, although our methods can be used for other basis functions. For example, denoting $\mathbf{B}(s)=[B_{1}(s),\ldots,B_{m}(s)]^{T}\in\mathbb{R}^{m}$ as a set of $m$ B-spline basis functions, we can represent the functional coefficients $\beta_{r}(s)=\sum_{d=1}^{m}\theta_{r,d}B_{d}(s)$. In practice, the dimension $m$ can differ across $r$. We denote $\bm{\theta}_{r}=[\theta_{r,1},\ldots,\theta_{r,m}]^{T}\in\mathbb{R}^{m}$ as an unknown parameter vector associated with covariate $r$, $\mathbf{B}=[\mathbf{B}(s_{1}),\ldots,\mathbf{B}(s_{L})]^{T}\in\mathbb{R}^{L\times m}$, and the linear predictor for a full observation of the functional outcome as $\bm{\eta}_{i,j}=[\eta_{i,j}(s_{1}),\ldots,\eta_{i,j}(s_{L})]^{T}=\mathbf{B}\bm{\theta}_{0}+\sum_{r=1}^{q}X_{i,j,r}\mathbf{B}\bm{\theta}_{r}$. We further define $\mathbb{X}_{i,j}=[\mathbf{B},~X_{i,j,1}\mathbf{B},\ldots,X_{i,j,q}\mathbf{B}]\in~\mathbb{R}^{L\times p}$, where $p=m(1+q)$. We then have that $\bm{\eta}_{i,j}=\mathbb{X}_{i,j}\bm{\theta}$, where $\bm{\theta}=[\bm{\theta}_{0}^{T},~\bm{\theta}_{1}^{T},~\ldots,~\bm{\theta}_{q}^{T}]^{T}\in\mathbb{R}^{p}$. Thus, we can estimate the functional coefficient vector, $\bm{\beta}_{r}=[\beta_{r}(s_{1}),~\ldots,~\beta_{r}(s_{L})]^{T}\in\mathbb{R}^{L}$, by estimating $\bm{\theta}$ and calculating $\widehat{\bm{\beta}}_{r}=\mathbf{B}\widehat{\bm{\theta}}_{r}$.
We semi-parametrically estimate the $\bm{\theta}$ with the penalized spline-based fGEE proposed in Chen et al. (2013). This assumes no likelihood and, if $\mu_{i,j}(s)$ is correctly specified, yields valid inference for $\{\bm{\beta}_{r}(s)\}_{s,r}$ even if $\text{Cov}(\bm{Y}_{i}\mid\mathbf{X}_{i})$ is misspecified. Specifically, the mean model parameters $\bm{\theta}$ are estimated as the root of the penalized estimating equation
@@
where $\mathbb{V}_{i}\in\mathbb{R}^{n_{i}L\times n_{i}L}$ is the working covariance matrix for cluster $i$ (whose true covariance matrix is $\mathbb{V}_{i}^{*}$), $\mathbb{D}_{i}=\frac{\partial\bm{\mu_{i}}(\bm{\theta})}{\partial\bm{\theta}}$, $\mathbb{X}_{i}=[\mathbb{X}_{i,1}^{T},\ldots,\mathbb{X}_{i,n_{i}}^{T}]^{T}\in\mathbb{R}^{n_{i}L\times p}$, and $\bm{\mu}_{i}=[\bm{\mu}_{i,1}^{T},\ldots,\bm{\mu}_{i,n_{i}}^{T}]^{T}\in\mathbb{R}^{n_{i}L}$. The pre-specified penalty matrix, $\mathbb{S}\in\mathbb{R}^{p\times p}$, is associated with the diagonal matrix of smoothing parameters $\Lambda\in\mathbb{R}^{p\times p}$ (see below for details). Although no likelihood is adopted, the estimating equation (2) can be derived from the score equations from, for example, an exponential dispersion family (Liang and Zeger, 1986); we add the penalty term for improved estimation in finite samples. Compared to a working independence matrix (e.g., $\mathbb{V}_{i}=\sigma^{2}I_{n_{i}L}$), estimation of $\bm{\theta}$ can be made more efficient and accurate by exploiting correlation, in functional and longitudinal directions, by choosing the working $\mathbb{V}_{i}$ to estimate $\mathbb{V}_{i}^{*}$. Although such choices for $\mathbb{V}_{i}$ in this fGEE model yield desirable statistical properties for longitudinal FDA, estimation is computationally intensive: estimating $\bm{\theta}$ based on equation (2) requires inversion of the $n_{i}L\times n_{i}L$ covariance matrix $\mathbb{V}_{i}$ for each cluster $i$, at each step in an optimization procedure.